Two small spheres separated 3cm apart carry equal charges. How many excess electrons must be placed on each sphere if the force of repulsion between them is to be 10^17N?

Grade:12th pass

1 Answers

Arun
4 years ago
 

let the charge on spheres is q

so F= 1/4∏ε qq/.032

  given F= 10-19

       1/4∏ε qq/.03=10-19 

 9 *109 * q2 /.03=10-19  

               q=10-16

1 electron has  1.6 * 10-19 C  

 ne =q

 n=q/e =10-16/1.6 * 10-19 C 

           =625